|Version 1 (modified by 11 months ago) ( diff ),|
APO 3.5-m Users Committee Telecon, 7/02/19
Attending: Mark Klaene (APO), Bill Ketzeback (APO), Russet McMillan (APO), Kevin Schlaufman (JHU), Ben Williams (UW), Kal Kadlec (UW), John Wisniewski (OU), Anne Verbiscer (UVa), Eric Bellm (UW), Joanne Hughes Clark (Seattle U), John Bally (CU), Sarah Tuttle (UW), Moire Prescott (NMSU), Nancy Chanover (NMSU)
User feedback and comments from institutional representatives:
- Colorado - nothing to report
- FGCU - no report
- Georgia State - no report
- JHU - nothing to report
- NAPG -no report
- NMSU - Question about what the plan is for KOSMOS time scheduled prior to September.
- Oklahoma -no report
- Seattle - no report
- UVa - nothing to report
- Washington - nothing to report
- Wyoming - no report
Telescope and Instruments Report:
Mark's detailed report is included below, followed by additional information discussed during today's call.
3.5-m Telescope and Instruments Highlights, 5/30/2019 - 6/26/2019
Reasonably good weather has led to some good observing. No moderate issues. We hosted the Strategic Planning meeting. Reminder to all users: summer shutdown is scheduled to run from July 15-24. We expect to be back on sky the night of July 19 for engineering. Based on the weather that could go quickly or could take longer than anticipated. Staff members are making plans and upgrades in preparation for the summer shutdown and the arrival of KOSMOS.
The telescope has been performing nominally.
DIS: The DIS blue camera continues to see contamination issues reappear rapidly after servicing while the red camera performance is deteriorating more slowly. Thus there are no plans to service DIS blue at this time. DIS red got a little worse but is still good.
TripleSpec is operational.
Agile is operational.
Echelle is operational. The new Ecam is in place and operable. Windowing is still in test mode.
NIC-FPS is operational.
ARCTIC is operational; we are working on some minor filter issues.
Operating nominally. From the Strategic Planning there was no desire for a complete makeover of ARCSAT due to cost. We will continue to band-aid and address issues as time and resources allow, however, we don't plan to do a major hardware or software upgrade.
Additional telescope and instrument discussion:
Summer shutdown starts on July 15 and will last for 5 days, plus a few days for engineering on-sky. Expect to be back for science on or before July 24. There are no future plans to continue servicing DIS Blue as the last half-dozen attempts were fruitless, with the contamination returning just a few days post-servicing. The new ecam was installed and is available. There remains a small issue with ecam windowing within TUI; this capability is currently broken but the site IT staff is working on a fix. Q3 passwords were sent out late but should be in mail today (7/2). If anyone needs passwords urgently, call the observatory.
There was a problem with some header cards not being populated correctly for data taken on all instruments. Namely, the UTC and LST time fields are missing, but the Date-Obs is correct and can be used to calculate the correct LST for the start of the observation. The problem has now been fixed, but will have affected any data taken from June 7-July 1. UC reps should let users know that the problem with headers has now been fixed, but it affected all data between June 7-July 1. Someone should send a note to 35m-general about this.
ARCSAT was a major topic of discussion at the strategic planning meeting held in June. There are no plans for a big overhaul at this time. We are working on improving guiding and addressed the SurveyCam shutter issue. Users should let site staff know about any issues they encounter, but note that fixes may take time given the lack of budget/manpower devoted to ARCSAT.
The interorder light in ARCES is higher than desired, although better than it was last year. This may not affect many programs, but users should keep an eye out, particularly if they are doing emission line science. UC reps should mention this echelle interorder light issue to users.
Q3 3.5m and ARCSAT scheduling:
The Q3 3.5m and ARCSAT schedules are out. There are 8 open half-nights on the 3.5m in Q3, including some dark time. There are also some open weeks in August and September for ARCSAT. We are eager to find takers, particularly for the open time in the next couple weeks. There are a number of class trips that have been scheduled or are being planned for the fall. Please let the staff know well before the Q4 proposal deadline about any class trips that are being planned to allow for coordination.
NA2 Baffle Project:
Work is proceeding, the stray light analysis is being done by outside firm. The target time frame for completion is Nov-Dec.
The UW TEG is making good progress on getting the instrument ready for commissioning on the 3.5m. Lots of parts are in the process of being made or are out for coating. Once everything is made by end of July, everything will be assembled. There will be about a month of testing, with the final pre-ship review likely Aug 12 (will send email) and shipping at the beginning of September. Commissioning will take place during September. Documentation on KOSMOS capabilities is still a work in progress.
Users with time scheduled before KOSMOS arrives in September should plan to use a back-up instrument.
Response to Call for White Papers for 2020 Decadal Survey
The white paper is due on July 10. Nancy sent out a link with a framework for the document. Based on discussions at the Strategic Planning meeting, we already have a pretty decent start on what it should contain, but it could use some figures and more text. If you have a few moments, please take a look, and if possible, contribute paragraph or figure. Remember to put your name on it if you make a contribution. Nancy will distribute the document more broadly to solicit input from those outside of ARC as well. UC reps should encourage people to take a look at document and contribute.
Discussion of 25th Anniversary / Strategic Planning meeting in 2019
Nancy sent out a complete set of minutes (26 pages) and asked people to look them over; she received no suggestions for edits/changes, so the minutes are now considered done. Nancy will post the minutes to the wiki. The consensus was to ask UC reps to send this link to users, but to not post it to the broader 35m-general list.
Nancy would like a final draft of a Strategic Plan ready by September, which would leave enough time to iterate on it with meeting attendees and various stakeholders and finalize it by the ARC Board of Governors meeting in November.
We are making progress on some of the near term action items that came out of the meeting, e.g., improving the observing proposal form, working with NMSU communications on an APOLLO story, getting larger narrow-band filters for ARCTIC. We will update the UC committee as we make progress in these areas.
Data Analysis Working Group (DAWG)
There was widespread agreement at the Strategic Planning Meeting that APO needs to invest in software tools, quicklook reduction, data archiving, and data analysis tools for APO instruments. A skeleton group was formed, and some email work is already happening. The near term actions are to identify a chair for the working group chair (any volunteers?) and finalizing the charge of the working group, keeping in mind that it can change as needed since the goals may be clarified once work gets underway. We discussed whether we want to seek broader participation in the group or whether we should keep the group small. There were various suggestions: we should have someone who is knowledgable about each instrument; we should reach out to the user community and let group membership be a bit more organic; we should actively recruit certain people with specific expertise; perhaps consider a committee of 5 who would work on refining the charge. Once that is done, then the actual working group can be much larger. We ended with the idea of having a core group guiding the process overall, with some subgroups working on different pieces.
Open action items from previous meetings:
- Ben will send out a link to KOSMOS info for interested prospective users. STATUS: OPEN. This is intended to be a page where prospective users could learn about the general capabilities of KOSMOS at a glance. Sarah and Kal have been worrying more about getting KOSMOS ready for commissioning, so this remains open. Sarah envisions this to be an entry point with capabilities that will be built up to a full manual that they can link to the APO web site. Ben will work with Kal to give him examples of what we're looking for.
- All UC reps should notify their users that if their Q3 proposals require DIS Blue for their science, they should contact Russet ASAP and let her know. STATUS: CLOSED.
- Nancy will post the final agenda and logistical info for the strategic planning meeting by the end of the day. STATUS: CLOSED.
New action items from this meeting:
- UC reps should let users know that the problem with headers has now been fixed, but it affected all data between June 7-July 1. Nancy will send a note to 35m-general about this.
- UC reps should mention as their users to be on the lookout for echelle interorder light issues, particularly if doing emission line science.
- UC reps should notify users of any OPEN time remaining. [Update: Aug 9B and Aug 15A are still available.]
- UC reps should encourage people to take a look at the Astro2020 white paper and contribute.
- Nancy will post the minutes from the Strategic Planning meeting to the users' wiki.
- Nancy will contact current DAWG members to finalize the organization of this working group.
All Other Business
Nancy sent a request to send her links to any APO-related papers published in 2018-2019, including submitted papers, dissertations, theses. As was discussed at the Strategic Planning meeting, the number and scientific impact of publications is one of the metrics of success of the observatory, so it is important to track this on an annual basis. Nancy needs to receive all inputs by July 31 in order to ensure that APO gets included in the full publication comparison by Dennis Crabtree.
The next Users Committee telecon will be on Monday August 5 at 10:30 am MDT.